OpenStudy (anonymous):

Write each equation in standard form. y+3=2/5(x+9) a. 2x – 5y = 33 c. y = 2/5x+3/5 b. 2x – 5y = –3 d. 2x + 5y = 3

OpenStudy (anonymous):

change each equation to the form y=mx+c where m and c are some numbers (constants) by isolating y on left hand side and simplifying the right hand side

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

he meant yolo

OpenStudy (anonymous):

ok I will do 1 for you y+3=2/5(x+9) y+3=2x/5+18/5 y=2x/5+18/5-3 y=2x/5+3/5